A capacitor of capacitance 8.0 μF is connected to a battery of emf 6.0 V through a resistance of 24 Ω. Find the current in the circuit (a) just after the connections are made and (b) one time constant after the connections are made.

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. C = 8 μF, E = 6V , R = 24 Ω a) l = V/R = 6/24 = 0.25A b) q = Q(1 – e^-t/RC) = (8 * 10^–6 * 6) [1 – c^–1] = 48 * 10^–6 * 0.63 = 3.024 * 10^–5 V = Q/C = 3.024 * 10^-5/8 * 10^-6 = 3.78 E = V + iR ⇒ 6 = 3.78 + i24 ⇒ i = 0.09 Å